# Assessing and improving reliability of neighbor embedding methods: a map-continuity perspective

### v1

@@ 2.2 Leave-one-out as a general diagnosis technique
where the approximation is based on the LOO assumption $\widetilde{{\bm{Y}}}\approx{\bm{Y}}$. This approximation allows us to decouple the dependence of $\widetilde{\bm{y}}_{i}$ on $\bm{x}$}.
Figure 2: Diagrams showing the idea of Leave-one-out (LOO) and LOO-map. a Idea of LOO. Adding one input data does not significantly change the overall positions of embedding points. The assumption allows us to analyze the properties of the embedding map over the entire input space via an approximated loss which we call LOO loss. b We introduce a global embedding map (LOO-map) ${\bm{f}}({\bm{x}})={\rm argmin}_{{\bm{y}}}L({\bm{y}};{\bm{x}})$ defined in the entire input space as an approximation to the neighbor embedding method $\mathcal{A}$.
@@
We empirically validate the LOO assumption by showing that $\bm{Y}$} and $\widetilde{\bm{Y}}$ are very close for large sample size $n$. Define the normalized error between embeddings before and after deleting a data point by
$\epsilon_{n}=\frac{1}{\|{\bm{Y}}\|_{F}}\|{\bm{Y}}-\widetilde{\bm{Y}}\|_{F}$ (4)
where $\|\cdot\|_{F}$ means the Frobenius norm of a matrix. A sufficiently small $\epsilon_{n}$ will support the approximation in our derivation of the surrogate map. In implementation, due to the highly non-convex nature of the loss function, the embedding $\bm{Y}$} is usually a local minimizer. We use this $\bm{Y}$} as the initialization for the second run of the optimization algorithm to compute the new embedding $\widetilde{{\bm{Y}}}$ (see Methods A.1).
We calculate this error extensively on both simulated and real datasets. The results support our LOO assumption (Table 1). We measured the approximation error $\epsilon_{n}$ across $20$ independent trials and with different sample size $n$. We observe that the approximation errors are small and generally decreasing in $n$, which aligns with our LOO assumption. Methods A.1 details the experiment setup and Supplementary Table S2 contains comprehensive results in support of our LOO assumption.
@@
Table 1: Empirical validation of LOO on both simulated datasets and real datasets. We measured the approximation error $\epsilon_{n}$ defined in Eqn. 4 across 20 independent trials and reported the average (and std) of $\epsilon_{n}$. We find that the errors are small and generally decreasing in $n$, which supports our LOO assumption.
2.3 LOO-map reveals intrinsic map discontinuities
@@
Figure 3: LOO loss landscape reveals the origins of two distortion patterns. a We illustrate two discontinuity patterns on simulated Gaussian mixture data. OI discontinuity: t-SNE embeds points into well-separated clusters and creates visual overconfidence. FI discontinuity: t-SNE with inappropriate perplexity creates many artificial fractures. b Origin of OI discontinuity: LOO loss contour plot shows distantly separated minima. We add a new input point $\bm{x}$} at one of the $4$ interpolated locations ${\bm{x}}=t{\bm{c}}_{1}+(1-t){\bm{c}}_{2}$ where $t\in\{0,0.47,0.48,1\}$ and then visualize the landscape of the LOO loss $L({\bm{y}};{\bm{x}})$ respectively by contour plots. Middle plots exhibit two minima (orange triangle) in well-separated valleys, thus causing a huge jump of the embedding point (as minimizer of the LOO loss) under a small perturbation of $\bm{x}$}. c Origin of FI discontinuity: We show LOO loss contour plots with interpolation coefficient $t\in\{0.2,0.4,0.6,0.8\}$. The plots show many local minima and irregular jumps. Under an inappropriate perplexity, loss landscape is consistently fractured. Numerous local minima cause an uneven trajectory of embedding points (dashed line) when adding $\bm{x}$} at evenly interpolated locations.
@@
By analyzing the LOO loss, we identify the two observed discontinuity patterns as a result of the map discontinuities of ${\bm{f}}({\bm{x}})$. We use t-SNE as an example to illustrate the main results.
@@
We generate mixture data by drawing $n=500$ input points from two overlapping 2D Gaussian distributions. We run the t-SNE algorithm twice under perplexity 5 and 50 respectively—which are two representative choices for the key hyperparameter of t-SNE. The visualization plots confirm the existence of the two observed discontinuity patterns (Figure 3a). For OI discontinuity, the mixed points are embedded to the boundaries of two well-separated clusters, thereby creating misleadingly tight cluster structures. For FI discontinuity, the embedding points form fragmented small pieces, leading to many sub-clusters. Similar discontinuity patterns are also common among other neighbor embedding methods (Supplementary Figure S1).
@@
We trace the origins of the observed discontinuity patterns by using the LOO loss function. To this end, we add a single point $\bm{x}$} at varying locations to the input data and track how $\bm{x}$} is mapped. By visualizing the landscape of the LOO loss function $L({\bm{y}};{\bm{x}})$ at four different inputs $\bm{x}$}, we provide snapshots of the LOO-map ${\bm{x}}\mapsto{\rm argmin}_{{\bm{y}}}L({\bm{y}};{\bm{x}})$. More specifically, we choose the centers ${\bm{c}}_{1},{\bm{c}}_{2}$ of the two Gaussian distributions and consider the interpolated input ${\bm{x}}(t)=t{\bm{c}}_{1}+(1-t){\bm{c}}_{2}$ where $t\in[0,1]$. Since ${\bm{x}}(t)$ is mapped to the minimizer of the LOO loss, the loss valley reveals the trajectory of the corresponding embedding point ${\bm{y}}(t)$ under varying coefficients $t$.
@@
We find that the observed OI discontinuity is caused by a discontinuity point of ${\bm{f}}({\bm{x}})$ in the midpoint of two mixtures. To demonstrate this, we visualize the LOO loss landscape and the embedding of the added point ${\bm{x}}(t)$ at four interpolated locations where $t\in\{0,0.47,0.48,1\}$. There are two clear well-separately valleys in the LOO loss landscape when $t\approx 0.5$ (Figure 3b). As a result, the embedding point ${\bm{y}}(t)$ jumps from one local minimum to the other minimum under a tiny change of $t$. A further gradient field analysis shows a hyperbolic geometry around the discontinuity point of ${\bm{f}}({\bm{x}})$; see Section 2.10 for a mathematical description of the loss landscape.
@@
We also find that the FI discontinuity is caused by numerous irregular local minima of $L({\bm{y}};{\bm{x}})$ under an inappropriate choice of perplexity. This conclusion is supported by the observation that the loss landscape of $L({\bm{y}};{\bm{x}})$ is consistently irregular and contains many local valleys under a small perplexity (Figure 3c). Moreover, varying the interpolation coefficient $t$ from 0 to 1 at a constant speed results in an uneven trajectory of the embedding point ${\bm{y}}(t)$. Because of many irregularities, the embedding points tend to get stuck at these local minima, thus forming spurious sub-clusters. In addition, we find that larger perplexity typically lessens the severity of FI discontinuity (Supplementary Figure S2, S3).
2.4 LOO-map induces diagnosis scores for capturing topological changes
@@
OI discontinuity and FI discontinuity reflect the properties of ${\bm{f}}({\bm{x}})$ at two different levels: OI discontinuity is relatively global while FI discontinuity relatively local. To quantify the degree of discontinuity, we introduce two point-wise scores (see Methods A.2 for detailed definitions): (i) Perturbation scores capture the severity of OI discontinuity, and (ii) singularity scores capture the severity of FI discontinuity. For computational efficiency, both scores obtained are based on modifying individual embedding points instead of adding a new point so that we maintain $n$ data points in total. This is justified by the LOO assumption that modifying/adding/deleting one point does not change embedding points much, allowing using the partial loss as LOO loss.
@@
Briefly speaking, we define the $i$-th perturbation score as the change of an embedding point ${\bm{y}}_{i}$ in the Euclidean distance under a perturbation of an input point ${\bm{x}}_{i}$ of moderate length. As the data distribution is not known a prior, we search the perturbation directions using the top principal directions of the data (Methods A.2.1).
@@
We define the $i$-th singularity score as the inverse of the smallest eigenvalue of a Hessian matrix that is related to the sensitivity of the embedding point ${\bm{y}}_{i}$ under infinitesimal perturbation. Our derivation (Supplementary file Section E.1) indicates that small eigenvalues can cause significant local discontinuity, while a singular Hessian matrix leads to the sharpest discontinuity. We find infinitesimal perturbation well suited for capturing the local property of FI discontinuity. Detailed expressions for the singularity scores of t-SNE, UMAP and LargeVis are provided in Supplementary Section E.2.
@@ 2.5 Simulation studies
Figure 4: Simulation studies demonstrate the effectiveness of proposed scores. a Perturbation scores identify unreliable embedding points that have reduced uncertainty. Input points from 5-component Gaussian mixture data form separated clusters in the embedding space. t-SNE reduces perceived uncertainty for input points in the overlapping region (left), as captured by the label-dependent measurements entropy difference (middle). Our perturbation scores can identify the same unreliable embedding points without label information (right). b-c Singularity scores reveal spurious sub-clusters on Gaussian mixture data (b) and Swiss roll data (c). At a low perplexity, t-SNE creates many spurious sub-clusters. Embedding points receiving high singular scores at random locations are an indication of such spurious structure.
@@
We implement our proposed point-wise scores for the t-SNE algorithm as an example and we believe that similar conclusions and results can be observed when applied to other neighbor embedding methods. We apply our diagnostic scores on two types of simulated datasets. (i) Gaussian mixture data (Methods A.4.1): we generate $K$ centers $({\bm{\mu}}_{k})_{k=1}^{K}$ in the two-dimensional space, and then draw $n=700$ input points with $K=5$ and unequal mixture probabilities, and $n=800$ input points with $K=8$ and equal mixture probabilities. (ii) Swiss roll data (Methods A.4.2): we randomly generate $n=800$ points from the Swiss-roll-shaped manifold in the three-dimensional space.
@@
We apply perturbation scores to the 5-component Gaussian mixture data, where the t-SNE algorithm unfaithfully disperses the overlapping regions of the input data due to OI discontinuity, resulting in misleadingly distinct cluster boundaries in the embedding (Figure 4a left). Without any utilizing label information, our perturbation scores can identify the unreliable embedding points whose levels of uncertainty have been deceptively reduced by t-SNE (Figure 4a right). Meanwhile, the entropy difference uses the ground-truth labels to calculate the reduced class entropy (Methods A.3.2) in the embedding space, thus providing an objective evaluation of the degree of confidence (Figure 4a middle). Our perturbation scores align very well with entropy difference.
@@
Next, we apply singularity scores to both the $8$-component Gaussian mixture data and the Swiss roll data under two choices of perplexity (Figure 4b-c). For each embedding plot, we color points according to ground-truth information, singularity scores, and dichotomized singularity scores (namely converting to binary values with a threshold). We find that the embedding structures are visually different: a small perplexity creates many spurious sub-clusters while a large perplexity faithfully preserves the cluster structure and manifold structure of respective datasets. Moreover, the distribution of the dichotomized scores are qualitatively different: at a low perplexity many high-scoring embedding points are randomly located, while at a high perplexity there are fewer high-scoring embedding points.
@@
Moreover, we provide further quantitative measurements of the clustering quality for the $8$-component Gaussian mixture data based on three commonly used indices, namely DB index [43], within-cluster distance ratio (Methods A.3.4), and Wilks’ $\Lambda$ [44]. All three indices (small values are better) indicate that t-SNE visualizations with less severe FI discontinuity, i.e., lower singularity scores, achieve better clustering quality. In particular, DB index decreases from $0.7470$ to $0.3821$, within-cluster distance ratio decreases from $0.3625$ to $0.1287$, and Wilks’ $\Lambda$ decreases from $0.2244$ to $0.0960$. To further study the change in clustering quality, we generate 6 simulated datasets with varying cluster structures and dimensions. Across all datasets, we consistently observe an improvement in the clustering quality after tuning perplexity using the singularity scores: the DB index decreases by approximately 50% for each dataset, the within-cluster distance ratio decreases by 65%–91%, and Wilks’ $\Lambda$ decreases by 57%–99% after selecting a suitably larger perplexity (Supplementary Table S3).
2.6 Use case 1: detecting out-of-distribution image data
@@
One common practical issue for statistical methods or machine learning algorithms is distribution shift, where the training dataset and test dataset have different distributions, often because they are collected at different sources [45, 46, 47]. These test data are called out-of-distribution (OOD) data.
@@
In this case study, we identify one rarely recognized pitfall of t-SNE visualization: OOD data may become harder to discern in t-SNE embeddings because they tend to be absorbed into other clusters. Our perturbation score is able to identify the dubious embedding points from OOD data.
@@
We consider a standard ResNet-18 model [32, 48] trained on the CIFAR-10 dataset [49] as mentioned in Section 2.1 . We use the model to extract features of CIFAR-10 test dataset as well as an OOD dataset known as DTD (describable textures dataset) [50]. Ideally, we hope that visualizing the features of test images and OOD images would reveal the distribution shift. However, the t-SNE embedding shows that a fraction of OOD features are absorbed into regular CIFAR-10 clusters that maintain tight cluster boundaries (Figure 5a). Without knowing the label information, one may mistakenly assume that the misplaced OOD embedding points belong to the regular and well-separated classes in CIFAR-10. We find that the OI discontinuity is the culprit of the embedding misplacement. Our inspection of the OOD features in the original space (before running the t-SNE algorithm) shows that misplaced OOD data points appear to have mixed membership, resembling both CIFAR-10 and OOD data—thus their cluster membership is in fact less certain than what t-SNE suggests.
@@
Our perturbation scores can successfully identify most of these misplaced OOD embedding points (Figure 5b, c, d). The areas under the ROC curves (AUC) are on average 0.75 for the three selected clusters. Additionally, examining label prediction probabilities through the neural network reveals that the entropies of these probabilities are significantly correlated with the perturbation scores, with correlations of 0.49, 0.58, and 0.64 for the clusters shown in Figure 5b-d. These findings suggest that perturbation scores are effective in detecting OOD data and can help safeguard against misinterpretation of t-SNE visualizations.
@@
Figure 5: Perturbation scores detect out-of-distribution (OOD) image data. a We use pretrained ResNet-18 model to extract features of CIFAR-10 images and, as out-of-distribution data, of DTD texture images. Then we visualize the features using t-SNE. A fraction of OOD embedding points are absorbed into clusters that represent CIFAR-10 image categories such as deer, truck, and automobile. b-d Perturbation scores can effectively identify misplaced out-of-distribution data points. The ROC curves show the proportion of OOD points correctly identified by the perturbation scores.
2.7 Use case 2: enhancing interpretation of single-cell data
@@
Figure 6: Singularity scores inform selection of the perplexity parameter. Comparative t-SNE embeddings and the corresponding singularity scores at two different perplexities in mouse embryonic cell differentiation data (a) and in mouse brain chromatin accessibility data (c). The tuning-parameter perplexity has a large impact on t-SNE visualization qualitatively. At small perplexity, there are many spurious sub-clusters. Embeddings with high singular scores appear in random locations, which indicate the presence of such spurious structures and severe FI discontinuity. Plots of degree of FI discontinuity and neighborhood preservation versus perplexity are shown for mouse embryonic cell differentiation data (b) and for mouse brain chromatin accessibility data (d). We recommend choosing perplexity no smaller than the elbow point, above which randomly located points with high singular scores disappear except in the peripheries of the clusters, and accordingly, the neighborhoods of most points are more faithfully embedded (better neighborhood preservation).
The first dataset we examined is a single-cell RNA-seq data of mouse embryonic stem cells (mESCs) undergoing differentiation [51]. This dataset contains the single-cell RNA sequencing data of 421 mouse embryonic stem cells collected at 5 sampling time points during differentiation. The second dataset is mouse brain single-cell ATAC-seq data [52]; this dataset records a collection of genome-wide chromatin accessibility profiles of 3618 mouse brain cells, summarized as the gene activity scores for each cell (Data Availability). These cells are labeled by 6 different cell type annotations. In our experiments, we vary a key tuning parameter in the t-SNE algorithm, namely perplexity. Briefly speaking, a higher perplexity means that input points have more “neighbors” in the calculation of similarity scores $v_{i,j}$.
Through analysis of both datasets, we find that a small perplexity tends to create spurious sub-clusters (Figure 6a, c). Our singular scores can provide informative insights into the spurious clusters even without the ground-truth labels, as summarized below.
@@
    (Distribution difference) Embedding points with large singularity scores tend to appear in random and scattered locations if the perplexity is too small. In contrast, under appropriate perplexity, embeddings with large singular scores are mostly in the periphery of clusters.

### v2

@@ Results
where the approximation is based on the LOO assumption $\widetilde{{\mathbf{Y}}}\approx{\mathbf{Y}}$. This approximation allows us to decouple the dependence of $\widetilde{\mathbf{y}}_{i}$ on $\mathbf{x}$}. We then define the LOO-map as ${\mathbf{f}}:{\mathbf{x}}\mapsto{\rm argmin}_{{\mathbf{y}}}L({\mathbf{y}};{\mathbf{x}})$.
Figure 2: Diagrams showing the idea of Leave-one-out (LOO) and LOO-map. a Idea of LOO. Adding one input point does not significantly change the overall positions of embedding points. The assumption allows us to analyze the properties of the embedding map over the entire input space via an approximated loss which we call LOO loss. b We introduce a global embedding map (LOO-map) ${\mathbf{f}}({\mathbf{x}})={\rm argmin}_{{\mathbf{y}}}L({\mathbf{y}};{\mathbf{x}})$ defined in the entire input space as an approximation to the neighbor embedding method $\mathcal{A}$.
@@
We empirically validate the LOO assumption by showing that $\mathbf{Y}$} and $\widetilde{\mathbf{Y}}$ are very close for a large sample size $n$. Define the normalized error between embeddings before and after deleting a data point by
$\epsilon_{n}=\frac{1}{\|{\mathbf{Y}}\|_{F}}\|{\mathbf{Y}}-\widetilde{\mathbf{Y}}\|_{F},$ (4)
where $\|\cdot\|_{F}$ means the Frobenius norm of a matrix. A sufficiently small $\epsilon_{n}$ will support the approximation in our derivation of the LOO-map.
We calculate this error extensively on both simulated and real datasets (Methods). The results support our LOO assumption (Table 1, Supplementary Table S2). We observe that the approximation errors are small and generally decreasing in $n$, which validates our LOO assumption.
@@
Table 1: Empirical validation of LOO on both simulated datasets and real datasets. We measured the approximation error $\epsilon_{n}$ defined in Equation 4 across 20 independent trials and reported the average (and the standard deviation) of $\epsilon_{n}$. We find that the errors are small and generally decreasing in $n$, which supports our LOO assumption.
LOO-map reveals intrinsic map discontinuities
@@
Figure 3: LOO loss landscape reveals the origins of two distortion patterns. a We illustrate two discontinuity patterns on simulated Gaussian mixture data. OI discontinuity: t-SNE embeds points into well-separated clusters and creates visual overconfidence. FI discontinuity: t-SNE with an inappropriate perplexity creates many artificial fractures. b Origin of OI discontinuity: LOO loss contour plot shows distantly separated minima. We add a new input point $\mathbf{x}$} at one of the $4$ interpolated locations ${\mathbf{x}}=t{\mathbf{c}}_{1}+(1-t){\mathbf{c}}_{2}$ where $t\in\{0,0.47,0.48,1\}$ and then visualize the landscape of the LOO loss $L({\mathbf{y}};{\mathbf{x}})$ using contour plots in the space of $\mathbf{y}$}. The middle two plots exhibit two well-separated minima (orange triangle), which cause a huge jump of the embedding point (as the minimizer of the LOO loss) under a small perturbation of $\mathbf{x}$}. c Origin of FI discontinuity: We show LOO loss contour plots with interpolation coefficient $t\in\{0.2,0.4,0.6,0.8\}$. The plots show many local minima and irregular jumps. Under an inappropriate perplexity, the loss landscape is consistently fractured. Numerous local minima cause an uneven trajectory of embedding points (dashed line) when adding $\mathbf{x}$} at evenly interpolated locations.
@@
By analyzing the LOO loss, we identify the two observed discontinuity patterns as a result of the map discontinuities of ${\mathbf{f}}({\mathbf{x}})$. We use t-SNE as an example to illustrate the main results.
@@
We generate mixture data by sampling 500 points from two overlapping 2D Gaussian distributions and run t-SNE with two representative choices of perplexity, 5 and 50. The resulting visualizations confirm the two discontinuity patterns (Fig. 3a). OI discontinuity pushes mixed points to cluster boundaries, creating overly tight structures, while FI discontinuity fragments embeddings into small pieces, leading to many sub-clusters. Similar discontinuity patterns are also common among other neighbor embedding methods (Supplementary Fig. S1).
@@
We trace the origins of the observed discontinuity patterns by the LOO loss function. To this end, we add a single point $\mathbf{x}$} at varying locations to the input data and track how $\mathbf{x}$} is mapped. By visualizing the landscape of the LOO loss $L({\mathbf{y}};{\mathbf{x}})$ at four different inputs $\mathbf{x}$}, we provide snapshots of the LOO-map ${\mathbf{x}}\mapsto{\rm argmin}_{{\mathbf{y}}}L({\mathbf{y}};{\mathbf{x}})$. More specifically, we choose the centers ${\mathbf{c}}_{1},{\mathbf{c}}_{2}$ of the two Gaussian distributions and consider the interpolated input ${\mathbf{x}}(t)=t{\mathbf{c}}_{1}+(1-t){\mathbf{c}}_{2}$, $t\in[0,1]$. Since ${\mathbf{x}}(t)$ is mapped to the LOO loss minimizer, tracking the loss minima reveals the trajectory of the corresponding embedding point ${\mathbf{y}}(t)$ under varying $t$.
@@
We find that the observed OI discontinuity is caused by a discontinuity point of ${\mathbf{f}}({\mathbf{x}})$ in the midpoint of two mixtures. To demonstrate this, we visualize the LOO loss landscape and the embedding of the added point ${\mathbf{x}}(t)$ at four interpolated locations where $t\in\{0,0.47,0.48,1\}$. There are two clearly well-separated minima in the LOO loss landscape when $t\approx 0.5$ (Fig. 3b). As a result, the embedding point ${\mathbf{y}}(t)$ jumps abruptly between local minima with a slight change in $t$. A further gradient field analysis shows a hyperbolic geometry around the discontinuity point of ${\mathbf{f}}({\mathbf{x}})$ (Fig. 8).
@@
We also find that the FI discontinuity is caused by numerous irregular local minima of $L({\mathbf{y}};{\mathbf{x}})$ under an inappropriate choice of perplexity. This conclusion is supported by the observation that the loss landscape of $L({\mathbf{y}};{\mathbf{x}})$ is consistently irregular and contains many local valleys under a small perplexity (Fig. 3c). Moreover, varying the interpolation coefficient $t$ from 0 to 1 at a constant speed results in an uneven trajectory of the embedding point ${\mathbf{y}}(t)$. Because of many irregularities, the embedding points tend to get stuck at these local minima, thus forming spurious sub-clusters. In addition, we find that larger perplexity typically lessens FI discontinuity (Supplementary Fig. S2, S3).
LOO-map motivates diagnostic scores for capturing topological changes
@@
OI discontinuity and FI discontinuity reflect the properties of ${\mathbf{f}}({\mathbf{x}})$ at different levels: OI discontinuity is relatively global, while FI discontinuity is relatively local. To quantify their severity respectively, we introduce two point-wise scores (Methods): (i) perturbation scores for OI discontinuity and (ii) singularity scores for FI discontinuity. For computational efficiency, both scores are based on modifying individual input points instead of adding a new point so that we maintain $n$ data points in total. This approach is justified by the LOO assumption, allowing using the partial loss as LOO loss.
@@
Briefly speaking, we define the perturbation score as the amount of change of an embedding point ${\mathbf{y}}_{i}$ under the perturbation of an input point ${\mathbf{x}}_{i}$ of a moderate length. As the data distribution is not known a priori, we search the perturbation directions using the top principal directions of the data (Methods).
@@
We define the singularity score as the inverse of the smallest eigenvalue of a Hessian matrix that represents the sensitivity of the embedding point ${\mathbf{y}}_{i}$ under infinitesimal perturbations. Our derivation (Supplementary Methods A.1) reveals that small eigenvalues can produce substantial local discontinuities, whereas a singular Hessian matrix leads to the most severe discontinuity. We find that infinitesimal perturbations are particularly effective for capturing the local characteristics of FI discontinuities. Detailed expressions for the singularity scores of t-SNE, UMAP and LargeVis are provided in Supplementary Methods A.2.
@@
Figure 4: Simulation studies demonstrate the effectiveness of proposed scores. a Perturbation scores identify unreliable embedding points that have reduced uncertainty. Input points from 5-component Gaussian mixture data form separated clusters in the embedding space. t-SNE reduces perceived uncertainty for input points in the overlapping region (left), as captured by the label-dependent measurements, namely the entropy difference (middle). Our perturbation scores can identify the same unreliable embedding points without label information (right). b-c Singularity scores reveal spurious sub-clusters on Gaussian mixture data (b) and Swiss roll data (c). At a low perplexity, t-SNE creates many spurious sub-clusters. Embedding points receiving high singular scores at random locations are an indication of such spurious structures.
@@
We implement our proposed point-wise scores for t-SNE as an example. We evaluate our diagnostic scores on two types of simulated datasets (Methods): (i) 2D Gaussian mixture data with 5 centers (unequal mixture probabilities, $n=700$) and 8 centers (equal probabilities, $n=800$), and (ii) Swiss roll data, where $n=800$ points are sampled from a 3D Swiss-roll manifold.
@@
We apply perturbation scores to the 5-component Gaussian mixture data, where t-SNE creates misleadingly distinct cluster boundaries (Fig. 4a left). Without label information, our scores identify unreliable points with deceptively low uncertainty (Fig. 4a right). Meanwhile, the entropy differences use the ground-truth labels to calculate the reduced class entropies (Methods) in the embedding space, thus providing an objective evaluation of the degree of confidence (Fig. 4a middle). Our perturbation scores are closely aligned with the entropy differences.
@@
Next, we apply singularity scores to the 8-component Gaussian mixture and Swiss roll data under two perplexity settings (Fig. 4b-c). Each embedding is colored by ground-truth labels, singularity scores, and dichotomized singularity scores (binary thresholding). The embeddings differ visually: a low perplexity creates spurious sub-clusters, while a high perplexity preserves cluster and manifold structures. Additionally, the distributions of dichotomized scores vary: a low perplexity results in more high scores at randomly scattered locations, whereas a high perplexity yields fewer high-scoring points.
@@
Moreover, we quantitatively assess the clustering quality for the 8-component Gaussian mixture data using three indices: DB index [45], within-cluster distance ratio (Methods), and Wilks’ $\Lambda$ [46]. All three indices (small values are better) indicate that t-SNE visualizations with less severe FI discontinuity, i.e., lower singularity scores, achieve better clustering quality, with the DB index dropping from 0.5982 to 0.3038, the within-cluster distance ratio from 0.0480 to 0.0024, and Wilks’ $\Lambda$ from 0.0028 to $9.0\times 10^{-6}$. To further study the change in clustering quality, we generate 6 simulated datasets with varying cluster structures and dimensions. Across all datasets, tuning perplexity using singularity scores consistently improves clustering quality, reducing the DB index by approximately 50%, the within-cluster distance ratio by 65%-91%, and Wilks’ $\Lambda$ by 57%-99% (Supplementary Table S3).
Use case 1: detecting out-of-distribution image data
@@
One common practical issue for statistical methods or machine learning algorithms is the distribution shift, where the training dataset and test dataset have different distributions, often because they are collected at different sources [47, 48, 49]. These test data are called out-of-distribution (OOD) data.
@@
In this case study, we identify one rarely recognized pitfall of t-SNE visualization: OOD data may become harder to discern in t-SNE embeddings because they tend to be absorbed into other clusters. Our perturbation score is able to identify the misplaced OOD embedding points.
@@
We use a standard ResNet-18 model [34] trained on the CIFAR-10 dataset [50] to extract features of its test dataset and an OOD dataset known as DTD (describable textures dataset) [51]. Ideally, visualization of the features of test images and OOD images would reveal the distribution shift. However, the t-SNE embedding shows that a fraction of OOD features are absorbed into compact, well-defined CIFAR-10 clusters (Fig. 5a). Without the label information, one may mistakenly assume that the misplaced OOD embedding points belong to the regular and well-separated classes in CIFAR-10. We find that the embedding misplacement results from OI discontinuity. Our inspection of the original feature space shows that the misplaced OOD data points appear to have mixed membership, resembling both CIFAR-10 and OOD data—thus their cluster membership is, in fact, less certain than what t-SNE suggests.
@@
Our perturbation scores can successfully identify most of these misplaced OOD embedding points (Fig. 5b-d). The areas under the ROC curves (AUROC) are on average 0.75 for the three selected clusters. Compared with other methods aiming for OOD detection, our perturbation score demonstrates superior performance, with kernel PCA [52] achieving an average AUROC of 0.698 and the one-class support vector machine [53] achieving an average AUROC of 0.410 (Supplementary Fig. S4, Methods). Additionally, we use the prediction probabilities given by the neural network to calculate the entropy of each point, and find that the entropies significantly correlate with the perturbation scores; specifically, the correlations are 0.49, 0.58, and 0.64 for the selected clusters. These findings suggest that perturbation scores are effective in detecting OOD data and can help safeguard against misinterpretation of t-SNE visualizations.
@@
Figure 5: Perturbation scores detect out-of-distribution (OOD) image data. a We use a pretrained ResNet-18 model to extract features of CIFAR-10 images and, as out-of-distribution data, of DTD texture images. Then we visualize the features using t-SNE with perplexity 100. A fraction of OOD embedding points are absorbed into clusters that represent CIFAR-10 image categories such as deer, truck, and automobile. b-d Perturbation scores can effectively identify misplaced out-of-distribution data points. The ROC curves show the proportion of OOD points correctly identified by the perturbation scores.
Use case 2: enhancing interpretation of single-cell data
@@
Figure 6: Singularity scores inform the selection of the perplexity parameter. Comparative t-SNE embeddings and the corresponding singularity scores at two different perplexities in mouse embryonic cell differentiation data (a) and in mouse mammary epithelial cell data (c). The perplexity as a tuning parameter has a large impact on t-SNE visualization qualitatively. At a small perplexity, there are many spurious sub-clusters. Embeddings with high singular scores appear in random locations, which indicates the presence of such spurious structures and severe FI discontinuity. Plots of the degree of FI discontinuity and neighborhood preservation versus perplexity are shown for mouse embryonic cell differentiation data (b) and for mouse mammary epithelial cell data (d). We recommend choosing a perplexity no smaller than the elbow point, as this ensures that randomly positioned points with high singularity scores largely disappear, remaining only at cluster peripheries. Consequently, the neighborhoods of most points are embedded more faithfully, resulting in better neighborhood preservation score.
The first dataset we examined is single-cell RNA-seq data from 421 mouse embryonic stem cells (mESCs) collected at 5 sampling time points during differentiation [54]. The second dataset is another single-cell RNA-seq data from 25,806 mouse mammary epithelial cells across 4 developmental stages [55]. We also include our analysis on a mid-sized mouse brain chromatin accessibility data in the Supplementary File. Through analysis of the datasets, we find that a small perplexity tends to create spurious sub-clusters (Fig. 6a, c, Supplementary Fig. S5a). Our singularity scores can provide informative insights into the spurious clusters even without the ground-truth labels, as summarized below.
@@
    (Distribution difference) Embedding points with large singularity scores tend to appear in random and scattered locations if the perplexity is too small. In contrast, under an appropriate perplexity, embedding points with large singular scores are mostly in the periphery of clusters.